Question 1(1 point) Use the chart to complete the following question. An economy produces only three goods bookshelves, laptops, and frying pans. It also consumes kiwi fruit, which it imports. Calculate the fixed weight GDP growth rate between years 1 and 2 using year O as the base year. Enter your answer as #########. For example, For example, 34.75% is entered as 0.35, 34.47% is entered 0.34 YearBookshelves Laptops Kiwi ans Quantity Price Quantit Price Quantity Price Quantity Price 30 $50 $60 $50 0 40 30 10 $1,000 20 $25 1000 $1 $201000$1 $50 1000 $1 12 $1,100 S500 30 30 2 30 Question 2 (1 point) Use the chart to complete the following question. An economy produces only three goods bookshelves, laptops, and frying pans. It also consumes kiwi fruit, which it imports. Calculate the fixed weight GDP growth rate between years 1 and 2 using year 1 as the base year. Enter your answer as #########. For example, For example, 34.75% is entered as 0.35, 34.47% is entered 0.34 YearBookshelves Laptops ans Quantity Price Quanti $50 $60 $50 Price Quantity Price Quantity Price 10 $1,000 12 $1,100 30 20 30 30 0 30 40 30 $25 1000 $1 $201000$1 $50 1000 $1 2 $500
Answer #1

(Question 1)

Fixed weight GDP keeps quantity consumed at base year values and imported goods are excluded, since GDP does not include imports.

Cost of basket, Year 0 ($) = 50 x 30 + 1,000 x 10 + 25 x 20 = 1,500 + 10,000 + 500 = 12,000

Cost of basket, Year 1 ($) = 60 x 30 + 1,100 x 10 + 20 x 20 = 1,800 + 11,000 + 400 = 13,200

Cost of basket, Year 2 ($) = 50 x 30 + 500 x 10 + 50 x 20 = 1,500 + 5,000 + 1,000 = 7,500

Therefore,

CPI, year 0 = 100 (Base year CPI is always 100)

CPI, year 1 = (Cost of basket, year 1 / Cost of basket, year 0) x 100 = ($13,200 / $12,000) x 100 = 110

CPI, year 2 = (Cost of basket, year 2 / Cost of basket, year 0) x 100 = ($7,500 / $12,000) x 100 = 62.5

% Change in GDP growth rate = % Change in CPI = (62.5 / 110) - 1 = 0.56 - 1 = -0.43

NOTE: As per Answering Policy, 1st question is answered.